# Scalable and consistent embedding of probability measures into Hilbert spaces via measure quantization

### v1

@@ Abstract
This paper is focused on statistical learning from data that come as probability measures. In this setting, popular approaches consist in embedding such data into a Hilbert space with either Linearized Optimal Transport or Kernel Mean Embedding. However, the cost of computing such embeddings prohibits their direct use in large-scale settings. We study two methods based on measure quantization for approximating input probability measures with discrete measures of small-support size. The first one is based on optimal quantization of each input measure, while the second one relies on mean-measure quantization. We study the consistency of such approximations, and its implication for scalable embeddings of probability measures into a Hilbert space at a low computational cost. We finally illustrate our findings with various numerical experiments.
@@ 1 Introduction
Machine Learning (ML) problems modeling data as a set of $N$ probability measures are classical in numerous applied fields such as signal and image processing, computer vision or computational biology, see the surveys [26, 32, 34, 24]. In particular, this framework includes distribution regression [37, 42, 6, 5, 31] that consists in predicting a scalar or label response from predictors that are probability measures, and the problem of Principal Component Analysis (PCA) of a set of probability measures [45, 41, 8, 10] for the purpose of dimension reduction from so-called distributional data [3]. Performing standard ML tasks on a set of probability measures is not straightforward as the algorithms are usually designed to handle $N$ points from an Euclidean space rather than $N$ distributions. However, as most of these methods rely on correlations through an inner-product, a popular approach is to embed such distributional data into a Hilbert space in which the whole machinery of machine learning methods can be easily applied. The two most commonly used embeddings are as follows: the first one is based on Linearized Optimal Transport (LOT) [45, 14, 33] and arises from the field of Optimal Transport (OT) [40, 36], by leveraging the Riemannian-like geometry of the space of probability measures endowed with the Wasserstein distance [2]. The second one, known as Kernel Mean Embedding (KME) [34], relies on the use of kernel methods to map probability measures into a Reproducing Kernel Hilbert Space (RKHS).
However, the computational cost and the storing of such embeddings prevent their use in large-scale settings. This is often the case when observing $N$ empirical measures on point clouds $X^{(i)}=(X^{(i)}_{1},\cdots,X^{(i)}_{m_{i}})\in(\mathbb{R}^{d})^{m_{i}},1\leq i\leq N$, with a large number $m_{i}$ of observations. Such datasets are frequently found in flow cytometry [30], where observations collected from $N$ patients represent a considerable amount of cells, each characterized by $d$ bio-markers. For these single-cell data, one usually encounters point clouds of thousands to millions of events (that is $m_{i}\geq 10^{5}$) living in a feature space of dimension $d$ larger than $10$. Using either the LOT embedding or the KME from such raw data becomes critical as these approaches suffer from high computational costs as soon as the number $m_{i}$ of points per clouds is larger than a few thousands.
Therefore, a relevant question is the following: given a set of $N$ probability measures with large support size, how to efficiently compute an embedding into a Hilbert space that is statistically consistent with the embedding derived directly from the raw data ?
@@ 1.1 Main contributions
In this paper, we consider the problem of embbeding a set of $d$-dimensional input probability measures $(\mu^{(i)})_{i=1}^{N}$ into a Hilbert space at a low computational cost. To that end, we propose to employ a preliminary $K$-quantization step that is either based on optimal quantization of each input measure $\mu^{(i)}$ or on the quantization of the mean measure $\bar{\mu}=\frac{1}{N}\sum_{i=1}^{n}\mu^{(i)}$ as used in [16] for single-cell data analysis. The aim of this $K$-quantization step is to approximate $(\mu^{(i)})_{i=1}^{N}$ by discrete measures $(\nu_{K}^{(i)})_{i=1}^{N}$ with supports of size $K$, with $K$ typically small. Using the theory of measure quantization [19, 35], we validate both quantization approaches by showing (Prop. 3.4):
$\mathcal{W}_{2}^{2}\left(\frac{1}{N}\sum_{i=1}^{N}\delta_{\mu^{(i)}},\frac{1}{N}\sum_{i=1}^{N}\delta_{\nu^{(i)}_{K}}\right)=\mathop{}\mathopen{}O\mathopen{}\left(K^{-2/d}\right),$ (1)
where $\mathcal{W}_{2}$ denotes the $2$-Wasserstein distance (4) on $\mathcal{P}(\mathcal{P}(\mathcal{X}))$, the set of probability measures over $\mathcal{P}(\mathcal{X})$, which is itself the set of probability measures with support included in a compact set $\mathcal{X}\subset\mathbb{R}^{d}$.
The asymptotic result (1) allows to show the convergence of numerous statistics computed from the $\nu^{(i)}_{K}$’s to corresponding quantities for the $\mu^{(i)}$’s as $K\to+\infty$. These statistics include the Wasserstein barycenter and the Gram matrix of the pairwise inner-products of the measures embedded to a Hilbert space using either LOT or KME. The latter is a standard quantity used in machine learning tasks (e.g. PCA).
Finally, the soundness and consistency of our method is illustrated with numerical experiments on synthetic and real datasets. We also show that the method based on mean-measure quantization has computational advantages over optimal quantization of each input measure while preserving satisfactory performances, which justifies its use in large scale settings.
@@ 1.2 Related works
In [12, 39], quantization is employed to embed a set of $N$ probability measures into a finite-dimensional Euclidean space through measure vectorization. More precisely, given $N$ input measures $\mu^{(i)}$, a quantization of the mean measure $\bar{\mu}=\frac{1}{N}\sum_{i=1}^{n}\mu^{(i)}$ by $K$ centers $x_{1},\ldots,x_{K}$ in $\mathbb{R}^{d}$ is first done. Then, they map each measure $\mu^{(i)}$ to $v^{(i)}=(v^{(i)}_{1},\cdots,v^{(i)}_{K})$ a vector of the convex space $\mathbb{R}_{+}^{K}$, where $v^{(i)}_{k}$ roughly represents the mass spread from the measure $\mu^{(i)}$ around the center $x_{k}$. Yet, this embedding does not take into account the relative positions of the $K$ centers, and consistency in the sense (1) is not shown as we propose in this paper by endowing the set of quantized measures $\nu^{(i)}_{K}$ with the Wasserstein distance (3).
@@
Finally, the benefits of a preliminary quantization step have been studied in [7] to improve the standard estimator of the OT cost between two probability measures based on the plug-in of their empirical counterpart. Still, the simultaneous quantization of $N$ probability measures for the purpose of constructing consistent and scalable embeddings has not been considered so far.
@@ 1.3 Organization of the paper
Section 2 presents OT, the embeddings LOT and KME and the quantization principle. In Section 3, we describe our two quantization methods of a set of $N$ probability measures, and analyze their theoretical properties. Section 4 reports the results of numerical experiments using synthetic and real data, and compare the computational cost of both methods. The paper ends with a conclusion in Section 5. All proofs are deferred to two technical Appendices A and B, and additional numerical experiments are given in Appendix C.
@@ 2 Background
$\min\limits_{\pi\in\Pi(\rho,\mu)}\int_{\mathcal{X}\times\mathcal{X}}\|x-y\|^{2}\mathrm{d}\pi(x,y),$ (2)
where $\Pi(\rho,\mu)$ is the set of probability measures (or transport plans) on $\mathcal{X}\times\mathcal{X}$ with marginals $\rho$ and $\mu$. For $\pi^{*}$ a minimizer of (2), the Wasserstein metric between $\rho$ and $\mu$ is
$W_{2}(\rho,\mu)=\Bigl{(}\int_{\mathcal{X}\times\mathcal{X}}\|x-y\|^{2}\mathrm{d}\pi^{*}(x,y)\Bigr{)}^{1/2}.$ (3)
@@
$\mathcal{W}_{2}(\mathbb{P},\mathbb{Q})=\Bigl{(}\min\limits_{\gamma\in\Gamma(\mathbb{P},\mathbb{Q})}\int_{\mathcal{P}(\mathcal{X})\times\mathcal{P}(\mathcal{X})}W_{2}^{2}(\rho,\mu)\mathrm{d}\gamma(\rho,\mu)\Bigr{)}^{1/2},$ (4)
@@
Given an absolutely continuous (a.c.) measure $\rho$, we first recall that Brenier’s theorem [9] states that the optimal transport plan $\pi^{*}$ in (2) is supported on the graph of a $\rho$-a.s. unique push-forward map [^1] We recall that the pushforward operator of a measure $\rho$ in $\mathbb{R}^{d}$ is defined as the measure $T_{\#}\rho$ such that for all Borelian $B\subset\mathbb{R}^{d},T_{\#}\rho=\rho\left(T^{-1}(B)\right)$. $T_{\rho}^{\mu}:\mathcal{X}\rightarrow\mathbb{R}^{d}$. In other words, $\pi^{*}=(\mathrm{id},T_{\rho}^{\mu})_{\#}\rho$ and
$W_{2}^{2}(\rho,\mu)=\int_{\mathcal{X}}\|x-T_{\rho}^{\mu}(x)\|^{2}\mathrm{d}\rho(x).$ (5)
The LOT embedding then consists in mapping a probability measure $\mu$ into the tangent space $\mathcal{T}_{{\rho}}=\{T_{\rho}^{\mu}-\mathrm{id}\}_{\mu\in\mathcal{P}(\mathcal{X})}$ at $\rho$, that is a subspace of the Hilbert space $L^{2}({\rho},\mathbb{R}^{d})=\{v:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}~{}|~{}\int_{\mathbb{R}^{d}}\|v\|^{2}\mathrm{d}{\rho}<\infty\}$, endowed with the weighted $L^{2}$ inner product $\langle v_{1},v_{2}\rangle_{L^{2}({\rho})}=\int_{\mathbb{R}^{d}}v_{1}(x)^{T}v_{2}(x)\mathrm{d}{\rho}(x)$.
Now, given a positive definite kernel function $k:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$ and associated RKHS $\mathcal{H}$, the KME of $\mu\in\mathcal{P}(\mathcal{X})$ is the embedding $\phi:\mathcal{P}(\mathcal{X})\rightarrow\mathcal{H}$ defined by $\phi(\mu)=\int_{\mathcal{X}}k(x,\cdot)\mathrm{d}\mathbb{\mu}(x)$. When the kernel $k$ is characteristic [34], the map $\phi$ is injective and one can define a metric on $\mathcal{P}(\mathcal{X})$ called Maximum Mean Discrepancy $\mathrm{MMD}(\rho,\mu)=\|\phi(\rho)-\phi(\mu)\|_{\mathcal{H}}.$
@@
We conclude this section with reminders on the theory of quantization [19, 35]. A $K$-points quantization of an arbitrary probability measure $\mu$ aims at approximating $\mu$ by solving [19][Lemma 3.4]:
$\min\limits_{a\in\Sigma_{K},X\in(\mathbb{R}^{d})^{K}}W_{2}^{2}\Bigl{(}\mu,\sum\limits_{k=1}^{K}a_{k}\delta_{x_{k}}\Bigr{)},$ (6)
where $\Sigma_{K}$ is the probability simplex in $\mathbb{R}^{K}$.
@@
If $\mu$ is a.c., it follows from [28][Proposition 2] that minimizers of (6) over $X\in(\mathbb{R}^{d})^{K}$ exist and belong to the set of pairwise distinct points
$F_{K}=\{X=(x_{1},\cdots,x_{K})\in(\mathbb{R}^{d})^{K}~{}|~{}x_{k}\neq x_{\ell},\mbox{ if }k\neq\ell\}.$
Given $X\in F_{K}$, it is well-known, see e.g. [25, 18], that the minimizer $a^{*}$ of $\min\limits_{a\in\Sigma_{K}}W_{2}^{2}\Bigl{(}\mu,\sum\limits_{k=1}^{K}a_{k}\delta_{x_{k}}\Bigr{)}$ is unique, and verifies $a_{k}^{*}=\mu(V_{x_{k}})$ where $(V_{x_{k}})_{k=1}^{K}$ is the set of Voronoï cells induced by $X$:
$V_{x_{k}}=\bigl{\{}y\in\mathbb{R}^{d}~{}|~{}\forall\ell\neq k,\|x_{k}-y\|^{2}\leq\|x_{\ell}-y\|^{2}\bigr{\}}.$ (7)
Thereby, the quantization problem (6) rewrites as:
$\min\limits_{X\in(\mathbb{R}^{d})^{K}}W_{2}^{2}\Bigl{(}\mu,\sum\limits_{k=1}^{K}\mu(V_{x_{k}})\delta_{x_{k}}\Bigr{)}.$ (8)
@@
If the measure $\mu$ is discrete, then the closed form solution in variable $a$ of the quantization problem (6) remains valid provided that the definition (7) of the Voronoï cells is slighted modified as follows
V~x1:=Vx1and
@@
so that $(\widetilde{V}_{x_{k}})_{k=1}^{K}$ form a partition of $\mathbb{R}^{d}$ [19][Chapter 1] for $X=(x_{1},\ldots,x_{K})\in F_{K}$. Then, for such a partitioning, it follows from Lemma A.1 in Appendix A that
$\min\limits_{a\in\Sigma_{K}}W_{2}^{2}\Bigl{(}\mu,\sum\limits_{k=1}^{K}a_{k}\delta_{x_{k}}\Bigr{)}=W_{2}^{2}\Bigl{(}\mu,\sum\limits_{k=1}^{K}\mu(\widetilde{V}_{x_{k}})\delta_{x_{k}}\Bigr{)}=\int_{\mathbb{R}^{d}}\min_{1\leq k\leq K}\{\|x_{k}-y\|^{2}\}\mathrm{d}\mu(y),$
for any probability measure $\mu$. As a consequence, and unless otherwise stated, the results of the paper hold for discrete probability measures with the choice (9) as a Voronoï partition, which corresponds to a chosen enumeration order of the elements of the set $X$. However, when $\mu$ is a discrete measure, it is necessary to require the cardinality of its support to be larger than $K$ so that the minimizers of (6) belong to $F_{K}$ [28][Proposition 2], and the sets of the Voronoï partition (9) are pairwise distinct.
Given a $K$-points quantization, that is a minimizer $X^{*}\in(\mathbb{R}^{d})^{K}$ of (6), the quantization error of the probability measure $\mu$ is defined as
$\varepsilon_{K}(\mu)=\int_{\mathbb{R}^{d}}\min_{1\leq k\leq K}\{\|x^{*}_{k}-y\|^{2}\}\mathrm{d}\mu(y).$ (10)
@@
3 Statistical properties of probability measures embeddings via measure quantization
Figure 1: Illustration of the quantization methods on two 1D probability measures (left) with $K=2$. (1) Quantization of each measure. (2a) Quantization of the mean measure. (2b) Computation of the weights for each measure. The vertical lines are located on the Dirac positions and weights of the quantized measures.
Throughout this section, the supports of the probability measures $(\mu^{(i)})_{i=1}^{N}$ are included in a compact set $\mathcal{X}\subset\mathbb{R}^{d}$. To reduce the computational costs of solving ML problems involving $N$ measures $(\mu^{(i)})_{i=1}^{N}$ with large supports, we propose two quantization methods, illustrated in Figure 1, to approximate the $\mu^{(i)}$’s with discrete measures supported on $K$ points.
Optimal quantization of each input measure.
A first natural approach is to approximate each $\mu^{(i)}$ by its optimal quantization
$\tilde{\nu}^{(i)}_{K}=\sum\limits_{k=1}^{K}\tilde{a}_{k}^{(i)}\delta_{x_{k}^{(i)}},$ (11)
where the weights $\tilde{a}^{(i)}=(\tilde{a}_{k}^{(i)})_{1\leq k\leq K}$ and locations $X^{(i)}=(x_{k}^{(i)})_{1\leq k\leq K}$ are minimizers of (6) for $\mu=\mu^{(i)}$.

### v3

@@ Abstract
This paper is focused on statistical learning from data that come as probability measures. In this setting, popular approaches consist in embedding such data into a Hilbert space with either Linearized Optimal Transport or Kernel Mean Embedding. However, the cost of computing such embeddings prohibits their direct use in large-scale settings. We study two methods based on measure quantization for approximating input probability measures with discrete measures of small-support size. The first one is based on optimal quantization of each input measure, while the second one relies on mean-measure quantization. We study the consistency of such approximations, and its implication for scalable embeddings of probability measures into a Hilbert space at a low computational cost. Finally, we illustrate our methods and compare them with existing approaches through various numerical experiments.
@@ 1 Introduction
Machine Learning (ML) techniques that model data as a set of N probability measures play a crucial role in various applied fields, including signal and image processing, computer vision, and computational biology [30, 36, 38, 28]. In particular, this framework includes distribution regression [41, 47, 6, 5, 35] for which the predictors are probability measures and the responses are scalar. Principal Component Analysis (PCA) of probability measures [52, 46, 8, 11] is also a key problem for dimensionality reduction of distributional data [3]. Performing these and other standard ML tasks on probability measures is challenging since most algorithms are designed to deal with $N$ points in a Euclidean space rather than $N$ probability distributions. However, as most machine learning methods for data analysis depend on the notion of inner-product, a common approach is to embed distributional data into a Hilbert space. To achieve this, there are two popular embeddings which stand out in the literature. The first one is the Linearized Optimal Transport (LOT) embedding [52, 17, 37], which arises from the theory of Optimal Transport (OT) [45, 40] and leverages the Riemannian-like geometry of the space of probability measures endowed with the Wasserstein distance [2]. The second one, known as Kernel Mean Embedding (KME) [38], relies on the use of kernel methods to map probability measures into a Reproducing Kernel Hilbert Space (RKHS).
However, the computational and storing costs of these embeddings make them impractical when dealing with probability measures with large support. This is often the case when observing $N$ empirical measures on point clouds $X^{(i)}=(X^{(i)}_{1},\cdots,X^{(i)}_{m_{i}})\in(\mathbb{R}^{d})^{m_{i}},1\leq i\leq N$, with a large number $m_{i}$ of observations. Such datasets are frequently found in flow cytometry [34], where observations collected from $N$ patients represent a considerable amount of cells, each characterized by $d$ bio-markers. For these single-cell data, one usually encounters point clouds of thousands to millions of events (that is $m_{i}\geq 10^{5}$) living in a feature space of dimension $d$ larger than $10$. Using either the LOT embedding or the KME from such raw data becomes critical as these approaches suffer from high computational costs as soon as the number $m_{i}$ of points per clouds is larger than a few thousands.
Given a set of $N$ probability measures with large support size, how to efficiently compute an embedding into a Hilbert space that is statistically consistent with the embedding derived directly from the raw data ?
@@ 1.1 Main contributions
In this paper, we consider the problem of embbeding a set of $d$-dimensional input probability measures $(\mu^{(i)})_{i=1}^{N}$ into a Hilbert space at a low computational cost. To that end, we propose to employ a preliminary $K$-quantization step that is either based on optimal quantization of each input measure $\mu^{(i)}$ or on the quantization of the mean measure $\bar{\mu}=\frac{1}{N}\sum_{i=1}^{n}\mu^{(i)}$ as used in [20] for single-cell data analysis. The aim of this $K$-quantization step is to approximate $(\mu^{(i)})_{i=1}^{N}$ by discrete measures $(\nu_{K}^{(i)})_{i=1}^{N}$ with supports of size $K$, with $K$ typically small. Using the theory of measure quantization [23, 39], we validate both quantization approaches by showing (see Theorem 3.4 below):
$\mathcal{W}_{2}^{2}\left(\frac{1}{N}\sum_{i=1}^{N}\delta_{\mu^{(i)}},\frac{1}{N}\sum_{i=1}^{N}\delta_{\nu^{(i)}_{K}}\right)=\mathop{}\mathopen{}O\mathopen{}\left(K^{-2/d}\right)\mbox{ as }K\to+\infty,$ (1.1)
where $\mathcal{W}_{2}$ denotes the $2$-Wasserstein distance (2.3) on $\mathcal{P}(\mathcal{P}(\mathcal{X}))$, the set of probability distributions over $\mathcal{P}(\mathcal{X})$, which is itself the set of probability measures with support included in a compact set $\mathcal{X}\subset\mathbb{R}^{d}$.
The asymptotic result (1.1) allows to show the convergence of numerous statistics computed from the $\nu^{(i)}_{K}$’s to corresponding quantities for the $\mu^{(i)}$’s as $K\to+\infty$. These include the Wasserstein barycenter and measures of statistical dispersion, from which clustering result can also be derived. In addition, we establish consistency results for quantities computed from the embeddings of the discrete measures $(\nu_{K}^{(i)})_{i=1}^{N}$ into a Hilbert space as $K$ increases, using either the LOT or KME framework. Precisely, we study the Gram matrix of pairwise inner-products of the embedded measures, which is a standard quantity used in machine learning applications. Within the LOT framework, we further prove the consistency of PCA computed from the embedded measures, corresponding to log-PCA [19] in the Wasserstein space.
Finally, the soundness and consistency of our method is illustrated with numerical experiments on synthetic and real datasets. We also show that the method based on mean-measure quantization has computational advantages over optimal quantization of each input measure while preserving satisfactory performances, which justifies its use in large scale settings. Finally, we compare our methods to classical embedding techniques for probability measures.
@@ 1.2 Related works
While quantization allows to approximate probability measures with a small set of points, other methods also aim at summarizing a dataset with representative samples. Within the framework of coresets [27, 15], one selects a subset of points such that solving a particular problem on this subset yields similar results than solving the problem on the entire dataset. In order to reduce the computational complexity of Gaussian Processes (GP) model, the principle of inducing variables, see e.g. [49], also allows for an approximation of the posterior by choosing a set of representative points and conditioning the GP on these points.
In [14, 44], quantization is employed to embed a set of $N$ probability measures into a finite-dimensional Euclidean space through measure vectorization. More precisely, given $N$ input measures $\mu^{(i)}$, a quantization of the mean measure $\bar{\mu}=\frac{1}{N}\sum_{i=1}^{n}\mu^{(i)}$ by $K$ centers $x_{1},\ldots,x_{K}$ in $\mathbb{R}^{d}$ is first done. Then, they map each measure $\mu^{(i)}$ to $v^{(i)}=(v^{(i)}_{1},\cdots,v^{(i)}_{K})$ a vector of the convex space $\mathbb{R}_{+}^{K}$, where $v^{(i)}_{k}$ roughly represents the mass of the the measure $\mu^{(i)}$ distributed around the center $x_{k}$. Yet, this embedding does not take into account the relative positions of the $K$ centers, and consistency in the sense (1.1) is not shown as we propose in this paper by endowing the set of quantized measures $\nu^{(i)}_{K}$ with the Wasserstein distance (2.2).
@@
Finally, the benefits of a preliminary quantization step have been studied in [7] to improve the standard plug-in estimator of the OT cost between two probability measures. Still, the simultaneous quantization of $N$ probability measures for the purpose of constructing consistent and scalable embeddings has not been considered so far.
@@ 1.3 Organization of the paper
Section 2 is devoted to some background on OT, the LOT and KME embeddings and the quantization principle. In Section 3, we describe our two quantization methods of a set of $N$ probability measures, and we prove their convergence in the sense of (1.1). In Section 4, we apply this result to derive the consistency of statistics derived from the quantized measures. In particular, we prove the consistency of various machine learning methods that take as inputs either the LOT embedding or KME of the discrete measures $(\nu_{K}^{(i)})_{i=1}^{N}$. Section 5 reports the results of numerical experiments using synthetic and real data, and the computational cost of both methods are discussed and compared. The paper ends with a conclusion in Section 6. All proofs are deferred to two technical Appendices A and B, and additional numerical experiments are given in Appendix C.
@@ 2 Background
$\min\limits_{\pi\in\Pi(\rho,\mu)}\int_{\mathcal{X}\times\mathcal{X}}\|x-y\|^{2}\mathrm{d}\pi(x,y),$ (2.1)
where $\Pi(\rho,\mu)$ is the set of probability measures (or transport plans) on $\mathcal{X}\times\mathcal{X}$ with marginals $\rho$ and $\mu$. For $\pi^{*}$ a minimizer of (2.1), the 2-Wasserstein metric between $\rho$ and $\mu$ is
$W_{2}(\rho,\mu)=\Bigl{(}\int_{\mathcal{X}\times\mathcal{X}}\|x-y\|^{2}\mathrm{d}\pi^{*}(x,y)\Bigr{)}^{1/2}.$ (2.2)
@@
$\mathcal{W}_{2}(\mathbb{P},\mathbb{Q})=\Bigl{(}\min\limits_{\gamma\in\Gamma(\mathbb{P},\mathbb{Q})}\int_{\mathcal{P}(\mathcal{X})\times\mathcal{P}(\mathcal{X})}W_{2}^{2}(\rho,\mu)\mathrm{d}\gamma(\rho,\mu)\Bigr{)}^{1/2},$ (2.3)
@@
Given an absolutely continuous (a.c.) measure $\rho$, we first recall that Brenier’s theorem [9] states that the optimal transport plan $\pi^{*}$ in (2.1) is supported on the graph of a $\rho$-a.s. unique push-forward map [^1] We recall that the pushforward of a measure $\rho$ in $\mathbb{R}^{d}$ by a measurable map $T$ is defined as the measure $T_{\#}\rho$ such that for all Borelian $B\subset\mathbb{R}^{d},T_{\#}\rho=\rho\left(T^{-1}(B)\right)$. $T_{\rho}^{\mu}:\mathcal{X}\rightarrow\mathbb{R}^{d}$. In other words, $\pi^{*}=(\mathrm{id},T_{\rho}^{\mu})_{\#}\rho$ and
$W_{2}^{2}(\rho,\mu)=\int_{\mathcal{X}}\|x-T_{\rho}^{\mu}(x)\|^{2}\mathrm{d}\rho(x).$ (2.4)
The LOT embedding then consists in mapping a probability measure $\mu$ to the function $T_{\rho}^{\mu}-\mathrm{id}$ which belongs to the Hilbert space $L^{2}({\rho},\mathbb{R}^{d})=\{v:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}~{}|~{}\int_{\mathbb{R}^{d}}\|v\|^{2}\mathrm{d}{\rho}<\infty\}$, endowed with the weighted $L^{2}$ inner product $\langle v_{1},v_{2}\rangle_{L^{2}({\rho})}=\int_{\mathbb{R}^{d}}v_{1}(x)^{T}v_{2}(x)\mathrm{d}{\rho}(x)$.
Now, given a positive definite kernel function $k:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$ and associated RKHS $\mathcal{H}$, the KME of $\mu\in\mathcal{P}(\mathcal{X})$ is the embedding $\phi:\mathcal{P}(\mathcal{X})\rightarrow\mathcal{H}$ defined by
$\phi(\mu)=\int_{\mathcal{X}}k(x,\cdot)\mathrm{d}\mathbb{\mu}(x).$ (2.5)
When the kernel $k$ is characteristic [38], the map $\phi$ is injective and one can define a metric on $\mathcal{P}(\mathcal{X})$ called Maximum Mean Discrepancy $\mathrm{MMD}(\rho,\mu)=\|\phi(\rho)-\phi(\mu)\|_{\mathcal{H}}.$
@@
We conclude this section with reminders on the theory of quantization [23, 39]. The optimal quantization of an arbitrary probability measure $\mu$ consists in approximating $\mu$ by a discrete measure that solves the following problem [23][Lemma 3.4]:
$\min\limits_{a\in\Sigma_{K},X\in(\mathbb{R}^{d})^{K}}W_{2}^{2}\Bigl{(}\mu,\sum\limits_{k=1}^{K}a_{k}\delta_{x_{k}}\Bigr{)},$ (2.6)
where $\Sigma_{K}$ is the probability simplex in $\mathbb{R}^{K}$ and $X=(x_{1},\cdots,x_{K})$, with $x_{k}\in\mathbb{R}^{d}$ for all $1\leq k\leq K$. For minimizers $a^{*}$ and $X^{*}$, a $K$-points quantization of $\mu$ is then defined by the discrete measure $\sum_{k=1}^{K}a^{*}_{k}\delta_{x^{*}_{k}}.$
@@
If $\mu$ is a.c., it follows from [32][Proposition 2] that minimizers of (2.6) over $X\in(\mathbb{R}^{d})^{K}$ exist and belong to the set of pairwise distinct points
$F_{K}=\{X=(x_{1},\cdots,x_{K})\in(\mathbb{R}^{d})^{K}~{}|~{}x_{k}\neq x_{\ell},\mbox{ if }k\neq\ell\}.$ (2.7)
Given a vector $X\in F_{K}$, it is well-known, see e.g. [29, 22], that the minimizer $a^{*}$ of $\min\limits_{a\in\Sigma_{K}}W_{2}^{2}\Bigl{(}\mu,\sum\limits_{k=1}^{K}a_{k}\delta_{x_{k}}\Bigr{)}$ is unique, and verifies $a_{k}^{*}=\mu(V_{x_{k}})$ where $(V_{x_{k}})_{k=1}^{K}$ is the set of Voronoï cells induced by $X$:
$V_{x_{k}}=\bigl{\{}y\in\mathbb{R}^{d}~{}|~{}\forall\ell\neq k,\|x_{k}-y\|^{2}\leq\|x_{\ell}-y\|^{2}\bigr{\}}.$ (2.8)
Thereby, the quantization problem (2.6) rewrites as:
$\min\limits_{X\in(\mathbb{R}^{d})^{K}}W_{2}^{2}\Bigl{(}\mu,\sum\limits_{k=1}^{K}\mu(V_{x_{k}})\delta_{x_{k}}\Bigr{)}.$ (2.9)
@@
If the measure $\mu$ is discrete, then the closed form solution in variable $a$ given in (2.9) remains valid provided that the definition (2.8) of the Voronoï cells is slightly modified as follows
Ux1:=Vx1and
@@
so that $(U_{x_{k}})_{k=1}^{K}$ form a partition of $\mathbb{R}^{d}$ [23][Chapter 1] for $X=(x_{1},\ldots,x_{K})\in F_{K}$. This modified definition is needed as some data points might end up on the boundaries of the Voronoï cells (2.8). In that case, the boundaries will have strictly positive measure and the quantization problem (2.6) can no longer be written in the closed-form solution (2.9) from Remark 2.1, that is when $\mu$ is a.c. For such a partitioning (2.10), it follows from Lemma A.1 in Appendix A that
$\min\limits_{a\in\Sigma_{K}}W_{2}^{2}\Bigl{(}\mu,\sum\limits_{k=1}^{K}a_{k}\delta_{x_{k}}\Bigr{)}=W_{2}^{2}\Bigl{(}\mu,\sum\limits_{k=1}^{K}\mu(U_{x_{k}})\delta_{x_{k}}\Bigr{)}=\int_{\mathbb{R}^{d}}\min_{1\leq k\leq K}\{\|x_{k}-y\|^{2}\}\mathrm{d}\mu(y),$
for any probability measure $\mu$. As a consequence, and unless otherwise stated, the results of the paper also hold for discrete probability measures with the choice (2.10) as a Voronoï partition, which corresponds to a chosen enumeration order of the elements of the vector $X$. However, when $\mu$ is a discrete measure, it is necessary to require the cardinality of its support to be larger than $K$ so that the minimizers of (2.6) belong to $F_{K}$ [32][Proposition 2], and the Voronoï cells in the partition (2.10) are pairwise distinct.
Given a $K$-points quantization, that is a minimizer $X^{*}\in(\mathbb{R}^{d})^{K}$ of (2.6), the quantization error of the probability measure $\mu$ is defined as
$\varepsilon_{K}(\mu)=\int_{\mathbb{R}^{d}}\min_{1\leq k\leq K}\{\|x^{*}_{k}-y\|^{2}\}\mathrm{d}\mu(y).$ (2.11)
@@
3 Consistency of measure quantizations
Figure 1: Illustration of the quantization methods on two (red and blue) 1D probability measures (left) with $K=2$. (1) Quantization of each measure. (2a) Quantization of the mean measure. (2b) Computation of the weights for each measure. The vertical lines represent the $K=2$ Dirac locations and weights for the quantized measures.
Throughout this section, the supports of the probability measures $(\mu^{(i)})_{i=1}^{N}$ are supposed to be included in a compact set $\mathcal{X}\subset\mathbb{R}^{d}$. To reduce the computational costs of solving ML problems involving $N$ measures $(\mu^{(i)})_{i=1}^{N}$ with large supports, we propose two quantization methods to approximate the $\mu^{(i)}$’s with discrete measures supported on $K$ points. These methods are described in the following and illustrated in Figure 1.
3.1 Two quantization schemes
Optimal quantization of each input measure
A first natural approach is to approximate each $\mu^{(i)}$ by its optimal quantization:
$\tilde{\nu}^{(i)}_{K}=\sum\limits_{k=1}^{K}\tilde{a}_{k}^{(i)}\delta_{\tilde{x}_{k}^{(i)}},$ (3.1)
where the weights $\tilde{a}^{(i)}=(\tilde{a}_{k}^{(i)})_{1\leq k\leq K}$ and locations $\widetilde{X}^{(i)}=(\tilde{x}_{k}^{(i)})_{1\leq k\leq K}$ are minimizers of (2.6) for $\mu=\mu^{(i)}$.